# Variational Signal Separation for Automotive Radar Interference Mitigation

### v1

@@ B-A Updates of Channel Weights
Using (21) to solve for the object channel weights $q_{\alpha}$, first note that the expectation of (46) with respect to $q_{\lambda}$ and the $q_{\gamma}$’s is obtained simply by applying $\lambda\rightarrow\hat{\lambda}$ and $\gamma\rightarrow\hat{\gamma}$. Neglecting constant terms independent of $\bm{\alpha}$, we are left with
@@ Appendix C Measurement of Interference
As introduced in Section IV-D, a measurement scheme whereby a separate interference signal is made available was devised. The parameters of the interference measurement can be found as Measurement II in Table I. The measurement took place in an indoor office with the victim and interfering sensors on a desk opposite each other. Crucially, the scenario was therefore completely static with no movement of objects. First, the location and duration of the interference bursts received by the victim radar were identified manually. The static object signal could then be estimated by taking the sample mean of the non-interfered data. This estimate was subtracted from the measured signal, approximately removing the object signal, leaving interference and measurement noise. Then, the time-limited interference bursts were cut out to exclude irrelevant noise-only sections.
In Section II, the effect of the AAF in the sensor hardware on the resulting interference burst was discussed. This is typically assumed to be known to some precision in the application by the algorithm designers of an automotive radar manufacturer. Section IV-B investigated the resilience of the proposed algorithm to a minor inaccuracy of this kind in the inference model. However, because the used hardware was a radar evaluation board from an external supplier, no details on the transfer function of the system was accessible to the authors. This can lead to a major model error. To mitigate this, the signal was additionally filtered digitally with a filter whose magnitude response dominates that of the sensor hardware. The concatenation of a high- and a low-pass was applied, both being 8-th order Butterworth filters with respective cut-off frequencies of $0.5\,\mathrm{MHz}$ and $5\,\mathrm{MHz}$. The known response of this digital filter chain was then used in our inference model, noting that this still constitutes a rather significant model error. Finally, the interference signal prepared in this way was added onto the measurement from Measurement I, yielding our test radar signal.

### v2

@@ B-A Updates of Channel Weights
Using (21) to solve for the object channel weights $q_{\alpha}$, first note that the expectation of (14) (i.e., the sum of (46) to (51)) with respect to $q_{\lambda}$ and the $q_{\gamma}$’s is obtained simply by applying $\lambda\rightarrow\hat{\lambda}$ and $\gamma\rightarrow\hat{\gamma}$. Neglecting constant terms independent of $\bm{\alpha}$, we are left with
@@ Appendix C Measurement of Interference
As introduced in Section IV-D, a measurement scheme whereby a separate interference signal is made available was devised. The parameters of the interference measurement can be found as Int. Measurement in Table I. The measurement took place in an indoor office with the victim and interfering sensors on a desk opposite each other. Crucially, the scenario was therefore completely static with no movement of objects. First, the location and duration of the interference bursts received by the victim radar were identified manually. The static object signal could then be estimated by taking the sample mean of the non-interfered data. This estimate was subtracted from the measured signal, approximately removing the object signal, leaving interference and measurement noise. Then, the time-limited interference bursts were cut out to exclude irrelevant noise-only sections.
In Section II, the effect of the AAF in the sensor hardware on the resulting interference burst was discussed. This is typically assumed to be known to some precision in the application by the algorithm designers of an automotive radar manufacturer. Section IV-B investigated the resilience of the proposed algorithm to a minor inaccuracy of this kind in the inference model. However, because the used hardware was a radar evaluation board from an external supplier, no details on the transfer function of the system was accessible to the authors. This can lead to a major model error. To mitigate this, the signal was additionally filtered digitally with a filter whose magnitude response dominates that of the sensor hardware. The concatenation of a high- and a low-pass was applied, both being 8-th order Butterworth filters with respective cut-off frequencies of $0.5\,\mathrm{MHz}$ and $5\,\mathrm{MHz}$. The known response of this digital filter chain was then used in our inference model, noting that this still constitutes a rather significant model error. Finally, the interference signal prepared in this way was added onto the measurement from Obj. Measurement, yielding our test radar signal.